The length of a rectangle is 6 inches longer than it is wide. If the area is 112 square inches, what are the dimensions of the rectangle?

The width, or shorter side is ___ inches

The length, or longer side is____ inches
Answer #1

Let width = x in hes

Therfore length = x+6 inch

Area = x(x+6)

112=x2+6x

x2+6x-112=0

x2+14x-8x-112=0

x(x+14)-8(x+14)+0

(x-8)(x+14)=0

x=8 inch

Therefore width = 8 inch

Length = 8+6 = 14 inch